OpenStudy (anonymous):

What values for q (0 < q < 2p) satisfy the equation? 2sqrt2 sin q +2= 0

OpenStudy (anonymous):

\[2\sqrt{2} \sin q +2 = 0 \]

OpenStudy (agent0smith):

q = 5pi/4, 7pi/4

OpenStudy (anonymous):

thanks :)